# Human and LLM Biases in Hate Speech Annotations: A Socio-Demographic Analysis of Annotators and Targets

### v3

@@ Introduction
Limitations and Future work. Some limitations must be acknowledged. Although extensive and rich, the social media data was collected from platforms with a marked US user-base and the annotations were obtained from US crowdsourcing contributors. Hence the data may primarily reflect US language, cultural norms, and perspectives. Moreover, the data is skewed in that some socio-demographic groups are over- or underrepresented and the data collection period was relatively short. These factors may limit the generalizability of the biases that we detected to other cultural contexts. Methodologically, our analysis of annotation bias, while informative, does not capture all dimensions of bias, such as contextual factors or annotators’ subjective experiences. Moreover, although we experimented with various prompts and LLM models, the specific prompting strategy and LLM versions used may have influenced the results. These limitations highlight the need for more representative datasets, improved methods for LLM personalization such as automatic prompt optimization (Schulhoff et al. 2024), and deeper investigations into specific biases in future research. Other than these, future work should also focus on intersectional studies, as combinations of attributes, rather than individual ones, can cause new biases or reinforce existing ones (Kim et al. 2020). To this end, the analyzed dataset appears particularly well-suited for comprehensive intersectional analyses given the availability of rich socio-demographic information for both annotators and targets.
@@ Acknowledgments
This work is partially supported by the European Union – Next Generation EU within the PRIN 2022 framework project PIANO (Personalized Interventions Against Online Toxicity), and by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab projects (Departments of Excellence).

### v4

@@ Introduction
Limitations and Future work. Some limitations must be acknowledged. While our methodology focuses on capturing relative differences in labeling behaviors between groups with and without specific attributes, we recognize that demographic imbalances in annotator and target populations may influence the observed patterns of bias. For example, certain annotator groups may be underrepresented, leading to a less comprehensive understanding of how their socio-demographic characteristics affect labeling decisions. Similarly, the overrepresentation of certain target groups as recipients of hate speech could amplify or obscure specific biases. Although our analysis reflects the data as collected, addressing these imbalances through stratified sampling or additional weighting strategies could further disentangle these effects in future work. Another possible limitation is that the social media data, although extensive and rich, was collected from platforms with a marked US user-base and the annotations were obtained from US crowdsourcing contributors. Hence the data may primarily reflect US language, cultural norms, and perspectives, potentially overlooking linguistic practices and cultural contexts from other regions or communities. This limitation may affect the applicability of the findings to other linguistic or geographic contexts. Future research should consider conducting similar experiments in different languages and cultural settings to enable meaningful comparisons with the results obtained in this study. Methodologically, our analysis of annotation bias, while informative, does not capture all dimensions of bias, such as contextual factors or annotators’ subjective experiences. Moreover, although we experimented with various prompts and LLM models, the specific prompting strategy and LLM versions used may have influenced the results. These limitations highlight the need for more representative datasets, improved methods for LLM personalization such as automatic prompt optimization (Schulhoff et al. 2024), and deeper investigations into specific biases in future research. Other than these, future work should also focus on intersectional studies, as combinations of attributes, rather than individual ones, can cause new biases or reinforce existing ones (Kim et al. 2020). To this end, the analyzed dataset appears particularly well-suited for comprehensive intersectional analyses given the availability of rich socio-demographic information for both annotators and targets.
@@ Acknowledgments
This work is partially supported by the European Union – Next Generation EU within the PRIN 2022 framework project PIANO (Personalized Interventions Against Online Toxicity); by the PNRR-M4C2 (PE00000013) “FAIR-Future Artificial Intelligence Research” - Spoke 1 ”Human-centered AI”, funded under Next Generation EU; and by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab projects (Departments of Excellence).
@@ Appendix A Appendix
Does this comment contain hate speech, defined as “bias-motivated, hostile and malicious language targeted at a person/group because of their actual or perceived innate characteristics, especially when the group is unnecessarily labeled (EXAMPLE: “jew bankers”).”
@@
If you can’t decide if the comment contains hate speech, you must label it as “maybe”